2.13 You are leasing a warehouse with annual lease payments of $100,000 $120,000, $135,000, $155,000, and $180,000 at the beginning of each year for the next five years respectively. If you earn 8 percent on your investment, how much in lump sum do you have to deposit today to be able to make all lease payments for the next five years?
Answer #1

Pv of 100000 at beginning of year 1 @ 8% = $100,000*1=$100,000

Pv of 120000 at beginning of year 2 @ 8% = $120,000*0.925925=$111,111

Pv of 135000 at beginning of year 3 @ 8% = $135,000*0.857338=$115,741

Pv of 155000 at beginning of year 4 @ 8% = $155,000*0.793832=$123,044

Pv of 180000 at beginning of year 5 @ 8% = $180,000*0.735030=$132,305

So the total investment required today is $582,201 which will earn 8% interest and we can pay all the lease payments of 5 years.